OpenStudy (anonymous):

What is the Sin (-25 pi/6) and Cos (-25 pi/3) Thanks

OpenStudy (***[isuru]***):

\[\sin (\frac{ -25\pi }{ 6 } ) = -\sin (\frac{ 25\pi }{ 6 } ) \] =\[-\sin (\frac{ 25\pi }{ 6 } ) = \sin( 4\pi\frac{ \pi }{ 6 })\] since 4pi = 2*2pi and it's multiplier of 2pi we just have to ignore that so....\[- \sin( 4\pi\frac{ \pi }{ 6 }) = \sin \frac{ \pi }{ 6 } \]

OpenStudy (***[isuru]***):

did ya get it.... or should I explain y I ignore 4pi out of the expression ?

OpenStudy (***[isuru]***):

can u work out Cos (-25 pi/3) ?

OpenStudy (anonymous):

so its gonna be -1/2 not 1/2 right? for the first one

OpenStudy (***[isuru]***):

OH GOD....am so sorry. I have frgt to put the negative sign ..... so it's should be -1/2

OpenStudy (***[isuru]***):

But the theory for the second one is different.... though \[\sin (-\theta) = - \sin \theta \] the cos value is different... for cos value it should be... \[\cos (- \theta) =\cos \theta\] *NOTHING CHANGES for cos value

OpenStudy (***[isuru]***):

can ya follow the same steps as for the sin value here... ? but remember wt I said earlier... just tell if ya want any more help! Hope this will help ya...!

OpenStudy (anonymous):

thank you so much got it

OpenStudy (***[isuru]***):

U r welcome!